# Scalable Bayesian Optimization via Focalized Sparse Gaussian Processes

### v1

@@ 5 Experiments
We compare FocalBO with representative sparse GP models used for Bayesian optimization, including SVGP (Hensman et al., 2013), WOGP (McIntire et al., 2016), and Vecchia GP (Jimenez and Katzfuss, 2023). We only run WOGP on synthetic functions due to its extremely low speed in dealing with the datasets in the remaining tasks. The number of inducing variables in sparse GP models is set as 50 for synthetic functions and as 200 for other tasks. The optimization performances are shown as mean $\pm$ 1 standard error for all considered problems over 10 independent trials.
@@ 5.1 Synthetic functions
We select Shekel and Michalewicz as the test functions, which are heterogeneous with both smooth and rigid regions. We also sample functions directly from Gaussian processes to evaluate algorithm performance under full BO assumption. For each function, we choose to use different acquisition functions to optimize: TS optimized by grid search, EI optimized by analytic gradient, and probability of improvement (PI) optimized by Monte Carlo gradient (Balandat et al., 2020). Optimization performances are shown in Figure 2. We observe that FocalBO significantly improves the performance of all acquisition functions compared to SVGP, and is able to consistently achieve top-tier performance over all problems. In Michalewicz function where a large fraction of the input space is flat, all baselines tend to increase the noise estimation to maintain a stationary prediction, while focalized GP is able to focus on the local search region and successfully optimize the function. Additional experiment with online samples as major data source is shown in Appendix B.2, where FocalBO still maintains comparable or better performance against baselines.
@@ 5.2 Robot morphology design
We compare FocalBO to several baselines over robot morphology design task from Design-Bench, which provides large offline dataset with an exact function oracle (Trabucco et al., 2022). The goal of the task is to optimize the morphological structure of D’Kitty robot (Ahn et al., 2020) to improve the simulation performance under RL controller. While the benchmark is initially designed for offline model-based optimization (MBO), it can also be used as an offline-to-online BO benchmark. In this task, we use the training dataset with 10,000 points and additionally evaluate 128 points on-the-fly with batch size of 4. EI is used as the base acquisition function for better optimizing with small batch size. We also try to combine FocalBO with TuRBO to optimize over the high-dimensional space, with the results shown in Figure 3. We observe that FocalBO achieves significant improvement from the initial data while other baselines struggle to obtain performance gain, even combined with TuRBO. FocalBO with TuRBO effectively extracts information from large offline dataset and is the first GP-based method to achieve top-tier performance reported by prior MBO works (Trabucco et al., 2021).
@@ 5.3 Human musculoskeletal system control
We further apply FocalBO to control a human arm musculoskeletal system (He et al., 2023) for the task of pouring liquid into a cup, as shown in Figure 4(a). To control the musculoskeletal system, we optimize a linear policy $\pi\in|A|\times|O|$, where $|A|=5$ and $|O|=117$ are the corresponding action and observation dimensions. The action dimension has been reduced from individual muscles to synergetic groups of muscles by applying principled component analysis to sampled action data from an RL agent (Appendix A.6). Although the original control dimension is reduced, the remaining 585-dimensional input space is still very high for existing high-dimensional BO algorithms. Therefore we consider a large offline-online setting, where we randomly sample 2000 points from the input space to serve as the offline dataset, and set the online budget as 3000 with batch size of 100. We use Thompson sampling as the base acquisition function. Figure 4(b) demonstrates that FocalBO outperforms other baselines, achieving higher maximum reward and faster convergence speed. Our supplementary video shows that the optimized policy is able to perform well on the task, demonstrating the successful application of FocalBO to high-dimensional control problems.
@@ A.3 GP training details
For all GP, we use Matern $\frac{5}{2}$ kernel with automatic relevance determination, and do not restrict the lengthscale or noise range. For each round of GP training, we fit GP hyperparameters (and variational parameters for focalized GP and SVGP) for 1000 epochs via Adam optimizer [Kingma and Ba, 2014] with learning rate as 0.01. For focalized GP and SVGP, we initialize the inducing points using Sobol sampler [Sobol’, 1967] over input space. all experiment are conducted on a server with Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz, NVIDIA-A100 and 512Gb memory.
@@ A.6 Human musculoskeletal system control
We use the musculoskeletal system from [He et al., 2023], which enables foward simlation with Mujoco [Todorov et al., 2012] and environment customization. We design the following reward for each environment step:
@@
We trained a Soft Actor-Critic (SAV) [Haarnoja et al., 2018] agent for 6M timesteps to collect task-related muscle activation data, and use principled component analysis to reduce the action dimension from 81 to 5.
@@ B.2 Optimization on synthetic functions with large online data
We choose Ackley and Hartmann, which are common-used test functions for BO community. We use the similar optimization setting in [Jimenez and Katzfuss, 2023]. The optimization performances are shown in Figure 9, where FocalBO is still able to achieve comparable or better performance when online samples dominates the data source.
@@ Appendix C Lemmas Used for Theoretical Implications of Focalized ELBO
(Corollary 19 in [Burt et al., 2020]). Let $k$ be a squared exponential kernel. Suppose that $N$ real-valued (onedimensional) covariates are observed, with identical Gaussian marginal distributions. Suppose the conditions of Theorem 13 are satisfied for some $R>0$. Fix any $\gamma\in(0,1]$. Then there exists an $M=\mathcal{O}(\log(N^{3}/\gamma))$ and an $\epsilon=\Theta(\gamma/N^{2})$ such if inducing points are distributed according to an $\epsilon$-approximate M-DPP with kernel matrix $K_{ff}$,
@@
(Proposition 1 in [Burt et al., 2020]). Suppose $2\text{KL}[Q\parallel P]\leq\gamma\leq\frac{1}{5}$. For any $x^{*}\in\mathcal{X}$, let $\mu_{1}$ denote the posterior mean of the variational approximation at $x^{*}$ and $\mu_{2}$ denote the mean of the exact posterior at $x^{*}$. Similarly, let $\sigma^{2}_{1},\sigma^{2}_{2}$ denote the variances of the approximate and exact posteriors at $x^{*}$. Then,
@@
(Assumption 4 in [Vakili et al., 2021]). (quality of the approximate prediction). For the approximate $\tilde{\mu}_{t}$, the exact $\mu_{t}$ and $\sigma_{t}$, and for all $x\in\mathcal{X}$,

### v2

@@ 5 Experiments
We compare FocalBO with representative sparse GP models used for Bayesian optimization, including SVGP (hensman2013gaussian), WOGP (mcintire2016sparse), and Vecchia GP (jimenez2023scalable). We only run WOGP on synthetic functions due to its extremely low speed in dealing with the datasets in the remaining tasks. The number of inducing variables in sparse GP models is set as 50 for synthetic functions and as 200 for other tasks. The optimization performances are shown as mean $\pm$ 1 standard error for all considered problems over 10 independent trials.
@@ 5.1 Synthetic functions
We select Shekel and Michalewicz as the test functions, which are heterogeneous with both smooth and rigid regions. We also sample functions directly from Gaussian processes to evaluate algorithm performance under full BO assumption. For each function, we choose to use different acquisition functions to optimize: TS optimized by grid search, EI optimized by analytic gradient, and probability of improvement (PI) optimized by Monte Carlo gradient (balandat2020botorch). Optimization performances are shown in Figure 2. We observe that FocalBO significantly improves the performance of all acquisition functions compared to SVGP, and is able to consistently achieve top-tier performance over all problems. In Michalewicz function where a large fraction of the input space is flat, all baselines tend to increase the noise estimation to maintain a stationary prediction, while focalized GP is able to focus on the local search region and successfully optimize the function. Additional experiment with online samples as major data source is shown in Appendix B.2, where FocalBO still maintains comparable or better performance against baselines.
@@ 5.2 Robot morphology design
We compare FocalBO to several baselines over robot morphology design task from Design-Bench, which provides large offline dataset with an exact function oracle (trabucco2022design). The goal of the task is to optimize the morphological structure of D’Kitty robot (ahn2020robel) to improve the simulation performance under RL controller. While the benchmark is initially designed for offline model-based optimization (MBO), it can also be used as an offline-to-online BO benchmark. In this task, we use the training dataset with 10,000 points and additionally evaluate 128 points on-the-fly with batch size of 4. EI is used as the base acquisition function for better optimizing with small batch size. We also try to combine FocalBO with TuRBO to optimize over the high-dimensional space, with the results shown in Figure 3. We observe that FocalBO achieves significant improvement from the initial data while other baselines struggle to obtain performance gain, even combined with TuRBO. FocalBO with TuRBO effectively extracts information from large offline dataset and is the first GP-based method to achieve top-tier performance reported by prior MBO works (trabucco2021conservative).
@@ 5.3 Human musculoskeletal system control
We further apply FocalBO to control a human arm musculoskeletal system (he2023self) for the task of pouring liquid into a cup, as shown in Figure 4(a). To control the musculoskeletal system, we optimize a linear policy $\pi\in|A|\times|O|$, where $|A|=5$ and $|O|=117$ are the corresponding action and observation dimensions. The action dimension has been reduced from individual muscles to synergetic groups of muscles by applying principled component analysis to sampled action data from an RL agent (Appendix A.6). Although the original control dimension is reduced, the remaining 585-dimensional input space is still very high for existing high-dimensional BO algorithms. Therefore we consider a large offline-online setting, where we randomly sample 2000 points from the input space to serve as the offline dataset, and set the online budget as 3000 with batch size of 100. We use Thompson sampling as the base acquisition function. Figure 4(b) demonstrates that FocalBO outperforms other baselines, achieving higher maximum reward and faster convergence speed. Our supplementary video shows that the optimized policy is able to perform well on the task, demonstrating the successful application of FocalBO to high-dimensional control problems.
@@ 6 Conclusion
Acknowledgments
This work is supported by STI 2030-Major Projects 2022ZD0209400. Correspondence to: Yanan Sui (ysui@tsinghua.edu.cn).
@@ A.3 GP training details
For all GP, we use Matern $\frac{5}{2}$ kernel with automatic relevance determination, and do not restrict the lengthscale or noise range. For each round of GP training, we fit GP hyperparameters (and variational parameters for focalized GP and SVGP) for 1000 epochs via Adam optimizer [kingma2014adam] with learning rate as 0.01. For focalized GP and SVGP, we initialize the inducing points using Sobol sampler [sobol1967distribution] over input space. all experiment are conducted on a server with Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz, NVIDIA-A100 and 512Gb memory.
@@ A.6 Human musculoskeletal system control
We use the musculoskeletal system from [he2023self], which enables foward simlation with Mujoco [todorov2012mujoco] and environment customization. We design the following reward for each environment step:
@@
We trained a Soft Actor-Critic (SAV) [haarnoja2018soft] agent for 6M timesteps to collect task-related muscle activation data, and use principled component analysis to reduce the action dimension from 81 to 5.
@@ B.2 Optimization on synthetic functions with large online data
We choose Ackley and Hartmann, which are common-used test functions for BO community. We use the similar optimization setting in [jimenez2023scalable]. The optimization performances are shown in Figure 9, where FocalBO is still able to achieve comparable or better performance when online samples dominates the data source.
@@ Appendix C Lemmas Used for Theoretical Implications of Focalized ELBO
(Corollary 19 in [burt2020convergence]). Let $k$ be a squared exponential kernel. Suppose that $N$ real-valued (onedimensional) covariates are observed, with identical Gaussian marginal distributions. Suppose the conditions of Theorem 13 are satisfied for some $R>0$. Fix any $\gamma\in(0,1]$. Then there exists an $M=\mathcal{O}(\log(N^{3}/\gamma))$ and an $\epsilon=\Theta(\gamma/N^{2})$ such if inducing points are distributed according to an $\epsilon$-approximate M-DPP with kernel matrix $K_{ff}$,
@@
(Proposition 1 in [burt2020convergence]). Suppose $2\text{KL}[Q\parallel P]\leq\gamma\leq\frac{1}{5}$. For any $x^{*}\in\mathcal{X}$, let $\mu_{1}$ denote the posterior mean of the variational approximation at $x^{*}$ and $\mu_{2}$ denote the mean of the exact posterior at $x^{*}$. Similarly, let $\sigma^{2}_{1},\sigma^{2}_{2}$ denote the variances of the approximate and exact posteriors at $x^{*}$. Then,
@@
(Assumption 4 in [vakili2021scalable]). (quality of the approximate prediction). For the approximate $\tilde{\mu}_{t}$, the exact $\mu_{t}$ and $\sigma_{t}$, and for all $x\in\mathcal{X}$,